OpenStudy (m):

what is ln of -e^-2.40/theta

OpenStudy (anonymous):

ln e^-x = 1/x

OpenStudy (m):

there is a negative in front of e

OpenStudy (m):

i don't think you can ln a negative

OpenStudy (m):

i'm guessing that negative doesn't matter in my case

OpenStudy (m):

i was hoping it would matter :)

OpenStudy (anonymous):

ln (-(1 / e^2.4) / theta), yea doesnt matter

OpenStudy (anonymous):

Log [-e^(-2.4)/θ] Log [-1/E^(2.4)*θ]] = Log[-0.09071795328941251/θ]

OpenStudy (m):

but you can't log a negative

OpenStudy (m):

-e^x/theta integrate from 2.40 to 0 and set it equal to .520 solve for theta

OpenStudy (m):

-e^2.40/theta - [-e^-0/theta] = .520 -e^-2.40/theta + 1 = .520 e^2.40/theta = .48 -2.40/theta = ln(.48) theta = 3.2699 i was trying to ln the limit 0 at first and that's why i was having problem. yay for perseverance :)